OpenStudy (anonymous):

Mara writes 120,000 in scientific notation and gets 1.2 × 10^5. A.Which of the following is a true statement? A.She is incorrect because the exponent should be 6. B.She is incorrect because the exponent should be negative. C.She is incorrect because 1.2 should be 0.12. D.She is correct.

jimthompson5910 (jim_thompson5910):

which one do you think it is?

OpenStudy (anonymous):

120,000 = 12 * 10^5 = 1.2 * 10^6

jimthompson5910 (jim_thompson5910):

Chlorophyll that is incorrect

jimthompson5910 (jim_thompson5910):

1.2 * 10^6 is 1.2 million and not 120 thousand

OpenStudy (anonymous):

Oop, wrong counting of "zero"

OpenStudy (anonymous):

120,000 = 12 * 10^4 = 1.2 * ....

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

12 * 10^4 = 1.2 * x Can you find x?

OpenStudy (anonymous):

umm idk